# Domain-Agnostic Stroke Lesion Segmentation Using Physics-Constrained Synthetic Data

### v1

@@ 1 Introduction
Segmenting brain pathologies in MRI is crucial for research and clinical applications but remains challenging due to the diversity of hospital imaging domains. Public datasets and benchmarks achieve high performance across pathologies but rely on consistent imaging sequences, such as T1-weighted (T1w), T2-weighted (T2w), and FLAIR, with similar parameters (e.g., echo time and flip angle). Even multi-site studies select data with overlapping sequences [22], limiting assessments to like-for-like comparisons: models trained on isotropic T1w are tested only on isotropic T1w, or co-registered T1w and T2w are tested together. In clinical settings, however, test domains rarely match training domains, creating a need for open-domain models capable of handling diverse MRI acquisitions.
Domain adaptation methods often assume specific target domains during training [10] or require many unlabelled target images at test time [26], neither of which suits clinical workflows, where only a single image may be available at test time. Domain-agnostic approaches, such as SynthSeg [3], which uses label-conditioned Gaussian Mixture Models (GMMs) to generate synthetic images with randomised tissue contrasts, have shown moderate success for lesion segmentation [2]. While effective for healthy tissue segmentation, SynthSeg struggles with pathologies like stroke, where lesions are heterogeneous and complex. Random sampling often produces unrealistic images, leading to poor generalisation, as demonstrated in [8], where models trained on synthetic data alone performed poorly in-domain. The lack of physical realism in synthetic images prevents models from learning true tissue relationships, causing errors on real clinical images.
@@
Despite its promise, qMRI acquisition is time-consuming, requiring multiple scans with specific sequences, which is impractical in time-critical settings like stroke. Consequently, qMRI data are scarce for training deep learning models. To overcome this, deep learning methods now estimate qMRI maps from standard sequences. Supervised approaches use real qMRI maps to synthesise images via forward models [5, 25], while unsupervised methods predict qMRI parameters from real images and validate them against input images via forward models [4]. These approaches enable qMRI generation without specialised acquisitions, making it feasible for synthetic data generation.
We propose two novel methods leveraging qMRI for domain-agnostic stroke lesion segmentation. The first, qATLAS, trains a qMRI estimation model to predict qMRI maps from MPRAGE images, commonly used in public datasets, enabling augmentation of the ATLAS dataset [18] with qMRI maps to simulate diverse MRI sequences while preserving acquisition physics. The second, qSynth, extends [8] by sampling synthetic qMRI maps from intensity priors derived from real qMRI data. Unlike methods that randomly sample intensities, qSynth ensures physical realism by generating qMRI maps and simulating sequences using forward models. This allows creation of large, physically plausible qMRI datasets for deep learning, improving stroke lesion segmentation on diverse imaging domains.
@@ 2 Methods
The qATLAS method involves training a qMRI estimation model to predict qMRI parameter maps - Proton Density (PD, arbitrary units), Longitudinal Relaxation Rate ($R_{1}$, $\text{s}^{-1}$), Effective Transverse Relaxation Rate ($R_{2}^{*}$, $\text{s}^{-1}$), and Magnetisation Transfer (MT, percentage units) - from MPRAGE images. We used a dataset of 51 subjects (22 healthy and 29 stroke patients), split into training and validation sets in an 80:20 ratio. The qMRI maps were estimated using the 3D-EPI method provided by the hMRI toolbox [23].
@@
Data Augmentation: To enhance the diversity of the training data, we applied several augmentation techniques using the MONAI library [7], including elastic and affine deformations, bias field augmentation, Gibbs ringing, Rician noise addition, and cropping to a size of $192\times 192\times 192$ voxels. Figure 1 illustrates examples of the augmented training data.
@@
Model Architecture and Training: We employed a U-Net model based on the nnUNet architecture [16] for the qMRI estimation task. The network consisted of five encoder stages, each containing two residual units with channel sizes of $(24,48,96,192,384)$, respectively. The model used GELU activations [14], instance normalisation [24], linear upsampling in the decoder, and a dropout rate of 0.1.
@@
The model was trained for a total of 200,000 iterations with a batch size of 1, using the AdamW optimiser [19] with a learning rate of $10^{-4}$ and parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay $\lambda=0.01$.. For the first 20,000 iterations, we used the L2 loss function. Subsequently, we employed a combined loss function inspired by [1], which included L1 and L2 losses on both the predictions and their spatial gradients:
@@
Additionally, we incorporated a perceptual loss based on the LPIPS metric [28], using a pre-trained Med3D model [9]. The perceptual loss $\text{PL}(\mathbf{y},\hat{\mathbf{y}})$ was weighted by a factor of 0.1 relative to the main reconstruction loss.
Figure 2 shows examples of the predicted qMRI parameter maps from input MPRAGE images in the ATLAS dataset [18].
@@
Generation of Synthetic qMRI Maps: We defined prior distributions for the qMRI parameters (PD, $R_{1}$, $R_{2}^{*}$, MT) for different tissue types (gray matter, white matter, cerebrospinal fluid, and lesions) based on the population used for the parameter estimation model in qATLAS. For each tissue, per-subject distribution parameters ($\mu$, $\beta$) were estimated, and then a population-level distribution parametrised by ($\mu_{\mu}$, $\mu_{\sigma}$, $\sigma_{\mu}$, $\sigma_{\sigma}$). For any tissue or lesion classes where real data is not available, it is feasible that this prior could be approximated from literature values. For each tissue label in the segmentation maps, we sampled the qMRI parameters from the corresponding priors to create synthetic qMRI maps. This method ensures that the synthesised qMRI maps accurately reflect the realistic range of tissue properties. All healthy tissue classes were acquired using the Multibrain SPM toolbox [6] to obtain classes of Gray Matter (GM), White Matter (WM), GM/WM partial volume, Cerebrospinal Fluid (CSF) and several non-brain tissues.
@@ 2.3 Physics-Based Generative Model
Noise Simulation: To simulate realistic MRI data, we added Rician noise to the generated images, modeling the noise characteristics of magnitude MRI data [12]. The noisy signal $S_{\text{noisy}}$ was computed as:
@@ 2.4 Segmentation Model Training
We trained nnUNet-based segmentation models [16] using both the qATLAS and qSynth datasets. Both methods are considered equally valid options for training domain-agnostic segmentation models. As baseline comparisons, we also train a model on the real MPRAGE images from the ATLAS dataset, as well as a standard synthetic data model using the public implementation of [8] (referred to here as ’Synth’). For both Synth and qSynth, we also show results for models trained with a mix of the synthetic data and the real ATLAS MPRAGE images.
Model Architecture and Training Details: The segmentation models were configured with PReLU activations [13] and one residual unit per block. For the qATLAS model, we predicted two classes: background and stroke lesion. For the qSynth model, trained with synthetic data generated from tissue labels, we predicted additional healthy tissue classes: gray matter (GM), white matter (WM), GM/WM partial volume, and cerebrospinal fluid (CSF).
Baseline and qATLAS models were trained using data from the ATLAS dataset (N=655, 419:105:131 train-val-test split), whilst Synth and qSynth models used the ATLAS lesion labels in combination with Multibrain healthy segmentation labels derived from the OASIS-3 dataset (N=2679, 2579:100 train-val split). Further details on the generative model used to create images from tissue labels can be found in the original implementation [8].
All models were optimised using a combined Dice and cross-entropy loss function, employing the AdamW optimiser [19] with a learning rate of $10^{-4}$ and parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay $\lambda=0.01$. A learning rate scheduler was applied as $\eta_{n}=\eta_{0}(1-\frac{n}{N})^{0.9}$, where $n$ is the current iteration and $N$ is the total number of iterations.
@@ 3 Experiments
Models were validated on several datasets. We evaluated models on the ATLAS hold-out test set (131 subjects, 1mm isotropic MPRAGE) for in-domain evaluation. For out-of-domain robustness, we used the ARC dataset (229 subjects with T2 images, 202 with T1, and 85 with FLAIR) [11, 17]. To further assess out-of-domain robustness in data closer resembling what is seen in a clinical workflow, we used a private dataset of 263 subjects with combinations of T2 and FLAIR images resliced to 2mm isotropic and spatially normalised to the MNI template, referred to here as PLORAS. The ISLES 2015 dataset (28 subjects, multi-modal MRI) [20] assessed model performance on acute lesions, highlighting physiological changes not covered by our augmentations.
At test-time, images are loaded, oriented to RAS, resliced to 1 mm slice thickness, histogram normalised and images z-scored to a zero mean and unit standard deviation. Sliding-window inference was performed using a patch size of $192\times 192\times 192$, using patches with a 50% overlap and a Gaussian weighting with $\sigma=0.125$ when combining patches. Test-time augmentation was additionally performed by averaging logits across all possible combinations of flips in the x, y and z axes. Predicted logits were then resliced to the original input image and a softmax applied to derive posterior probabilities. An argmax was then used to derive a binary lesion map. TTA is a popular heuristic method for increasing model robustness at test-time [27].
@@ 5 Conclusion
Looking ahead, combining qSynth with domain adaptation techniques could establish a robust foundation model adaptable across modalities. Incorporating multiple simulated sequences using strategies like weakly supervised learning [21] may further broaden applicability, particularly for pathologies commonly imaged with specific modality combinations. Additionally, advances in deep generative modelling could refine the qMRI generator, while translating multi-modal data into qMRI maps could improve the representation of physiological variability.
@@
An important extension is to explore whether the benefits of qMRI-based synthesis extend to other domains where synthetic data has been impactful, such as healthy brain parcellation [3], and super-resolution [15].
@@
6 Acknowledgement
LC is supported by the EPSRC-funded UCL Centre for Doctoral Training in Intelligent, Integrated Imaging in Healthcare (i4health) (EP/S021930/1), and the Wellcome Trust (203147/Z/16/Z and 205103/Z/16/Z). This research was supported by NVIDIA and used NVIDIA RTX A6000 48GB.

### v2

@@ 1 Introduction
Segmenting brain pathologies in MRI is crucial for research and clinical applications but remains challenging due to the diversity of hospital imaging domains. Public datasets and benchmarks achieve high performance across pathologies but rely on consistent imaging sequences, such as T1-weighted (T1w), T2-weighted (T2w), and FLAIR, with similar parameters (e.g., echo time and flip angle). Even multi-site studies select data with overlapping sequences [delarosa2024robustensemblealgorithmischemic], limiting assessments to like-for-like comparisons: models trained on isotropic T1w are tested only on isotropic T1w, or co-registered T1w and T2w are tested together. In clinical settings, however, test domains rarely match training domains, creating a need for open-domain models capable of handling diverse MRI acquisitions.
Domain adaptation methods often assume specific target domains during training [Dorent2023] or require many unlabelled target images at test time [wang2021tent], neither of which suits clinical workflows, where only a single image may be available at test time. Domain-agnostic approaches, such as SynthSeg [Billot2023], which uses label-conditioned Gaussian Mixture Models (GMMs) to generate synthetic images with randomised tissue contrasts, have shown moderate success for lesion segmentation [Billot2021]. While effective for healthy tissue segmentation, SynthSeg struggles with pathologies like stroke, where lesions are heterogeneous and complex. Random sampling often produces unrealistic images, leading to poor generalisation, as demonstrated in [chalcroft2024syntheticdatarobuststroke], where models trained on synthetic data alone performed poorly in-domain. The lack of physical realism in synthetic images prevents models from learning true tissue relationships, causing errors on real clinical images.
@@
Despite its promise, qMRI acquisition is time-consuming, requiring multiple scans with specific sequences, which is impractical in time-critical settings like stroke. Consequently, qMRI data are scarce for training deep learning models. To overcome this, deep learning methods now estimate qMRI maps from standard sequences. Supervised approaches use real qMRI maps to synthesise images via forward models [borges2021acquisitioninvariant, varadarajan2021unsupervised], while unsupervised methods predict qMRI parameters from real images and validate them against input images via forward models [Borges2023]. These approaches enable qMRI generation without specialised acquisitions, making it feasible for synthetic data generation.
We propose two novel methods leveraging qMRI for domain-agnostic stroke lesion segmentation. The first, qATLAS, trains a qMRI estimation model to predict qMRI maps from MPRAGE images, commonly used in public datasets, enabling augmentation of the ATLAS dataset [Liew2022] with qMRI maps to simulate diverse MRI sequences while preserving acquisition physics. The second, qSynth, extends [chalcroft2024syntheticdatarobuststroke] by sampling synthetic qMRI maps from intensity priors derived from real qMRI data. Unlike methods that randomly sample intensities, qSynth ensures physical realism by generating qMRI maps and simulating sequences using forward models. This allows creation of large, physically plausible qMRI datasets for deep learning, improving stroke lesion segmentation on diverse imaging domains.
@@ 2 Methods
The qATLAS method involves training a qMRI estimation model to predict qMRI parameter maps - Proton Density (PD, arbitrary units), Longitudinal Relaxation Rate ($R_{1}$, $\text{s}^{-1}$), Effective Transverse Relaxation Rate ($R_{2}^{*}$, $\text{s}^{-1}$), and Magnetisation Transfer (MT, percentage units) - from MPRAGE images. We used a dataset of 51 subjects (22 healthy and 29 stroke patients), split into training and validation sets in an 80:20 ratio. The qMRI maps were estimated using the 3D-EPI method provided by the hMRI toolbox [Tabelow2019].
@@
Data Augmentation: To enhance the diversity of the training data, we applied several augmentation techniques using the MONAI library [cardoso2022monai], including elastic and affine deformations, bias field augmentation, Gibbs ringing, Rician noise addition, and cropping to a size of $192\times 192\times 192$ voxels. Figure 1 illustrates examples of the augmented training data.
@@
Model Architecture and Training: We employed a U-Net model based on the nnUNet architecture [Isensee2020] for the qMRI estimation task. The network consisted of five encoder stages, each containing two residual units with channel sizes of $(24,48,96,192,384)$, respectively. The model used GELU activations [hendrycks2023gaussian], instance normalisation [ulyanov2017instance], linear upsampling in the decoder, and a dropout rate of 0.1.
@@
The model was trained for a total of 200,000 iterations with a batch size of 1, using the AdamW optimiser [loshchilov2019decoupled] with a learning rate of $10^{-4}$ and parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay $\lambda=0.01$.. For the first 20,000 iterations, we used the L2 loss function. Subsequently, we employed a combined loss function inspired by [pmlr-v102-baur19a], which included L1 and L2 losses on both the predictions and their spatial gradients:
@@
Additionally, we incorporated a perceptual loss based on the LPIPS metric [zhang2018unreasonable], using a pre-trained Med3D model [chen2019med3d]. The perceptual loss $\text{PL}(\mathbf{y},\hat{\mathbf{y}})$ was weighted by a factor of 0.1 relative to the main reconstruction loss.
Figure 2 shows examples of the predicted qMRI parameter maps from input MPRAGE images in the ATLAS dataset [Liew2022].
@@
Generation of Synthetic qMRI Maps: We defined prior distributions for the qMRI parameters (PD, $R_{1}$, $R_{2}^{*}$, MT) for different tissue types (gray matter, white matter, cerebrospinal fluid, and lesions) based on the population used for the parameter estimation model in qATLAS. For each tissue, per-subject distribution parameters ($\mu$, $\beta$) were estimated, and then a population-level distribution parametrised by ($\mu_{\mu}$, $\mu_{\sigma}$, $\sigma_{\mu}$, $\sigma_{\sigma}$). For any tissue or lesion classes where real data is not available, it is feasible that this prior could be approximated from literature values. For each tissue label in the segmentation maps, we sampled the qMRI parameters from the corresponding priors to create synthetic qMRI maps. This method ensures that the synthesised qMRI maps accurately reflect the realistic range of tissue properties. All healthy tissue classes were acquired using the Multibrain SPM toolbox [Brudfors_2020] to obtain classes of Gray Matter (GM), White Matter (WM), GM/WM partial volume, Cerebrospinal Fluid (CSF) and several non-brain tissues.
@@ 2.3 Physics-Based Generative Model
Noise Simulation: To simulate realistic MRI data, we added Rician noise to the generated images, modeling the noise characteristics of magnitude MRI data [Gudbjartsson1995]. The noisy signal $S_{\text{noisy}}$ was computed as:
@@ 2.4 Segmentation Model Training
We trained nnUNet-based segmentation models [Isensee2020] using both the qATLAS and qSynth datasets. Both methods are considered equally valid options for training domain-agnostic segmentation models. As baseline comparisons, we also train a model on the real MPRAGE images from the ATLAS dataset, as well as a standard synthetic data model using the public implementation of [chalcroft2024syntheticdatarobuststroke] (referred to here as ’Synth’). For both Synth and qSynth, we also show results for models trained with a mix of the synthetic data and the real ATLAS MPRAGE images.
Model Architecture and Training Details: The segmentation models were configured with PReLU activations [he2015delving] and one residual unit per block. For the qATLAS model, we predicted two classes: background and stroke lesion. For the qSynth model, trained with synthetic data generated from tissue labels, we predicted additional healthy tissue classes: gray matter (GM), white matter (WM), GM/WM partial volume, and cerebrospinal fluid (CSF).
Baseline and qATLAS models were trained using data from the ATLAS dataset (N=655, 419:105:131 train-val-test split), whilst Synth and qSynth models used the ATLAS lesion labels in combination with Multibrain healthy segmentation labels derived from the OASIS-3 dataset (N=2679, 2579:100 train-val split). Further details on the generative model used to create images from tissue labels can be found in the original implementation [chalcroft2024syntheticdatarobuststroke].
All models were optimised using a combined Dice and cross-entropy loss function, employing the AdamW optimiser [loshchilov2019decoupled] with a learning rate of $10^{-4}$ and parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay $\lambda=0.01$. A learning rate scheduler was applied as $\eta_{n}=\eta_{0}(1-\frac{n}{N})^{0.9}$, where $n$ is the current iteration and $N$ is the total number of iterations.
@@ 3 Experiments
Models were validated on several datasets. We evaluated models on the ATLAS hold-out test set (131 subjects, 1mm isotropic MPRAGE) for in-domain evaluation. For out-of-domain robustness, we used the ARC dataset (229 subjects with T2 images, 202 with T1, and 85 with FLAIR) [Gibson2024TheRepository, Johnson2024ProgressiveStroke]. To further assess out-of-domain robustness in data closer resembling what is seen in a clinical workflow, we used a private dataset of 263 subjects with combinations of T2 and FLAIR images resliced to 2mm isotropic and spatially normalised to the MNI template, referred to here as PLORAS. The ISLES 2015 dataset (28 subjects, multi-modal MRI) [Maier2017] assessed model performance on acute lesions, highlighting physiological changes not covered by our augmentations.
At test-time, images are loaded, oriented to RAS, resliced to 1 mm slice thickness, histogram normalised and images z-scored to a zero mean and unit standard deviation. Sliding-window inference was performed using a patch size of $192\times 192\times 192$, using patches with a 50% overlap and a Gaussian weighting with $\sigma=0.125$ when combining patches. Test-time augmentation was additionally performed by averaging logits across all possible combinations of flips in the x, y and z axes. Predicted logits were then resliced to the original input image and a softmax applied to derive posterior probabilities. An argmax was then used to derive a binary lesion map. TTA is a popular heuristic method for increasing model robustness at test-time [Wang_2019].
@@ 5 Conclusion
Looking ahead, combining qSynth with domain adaptation techniques could establish a robust foundation model adaptable across modalities. Incorporating multiple simulated sequences using strategies like weakly supervised learning [patel2022weakly] may further broaden applicability, particularly for pathologies commonly imaged with specific modality combinations. Additionally, advances in deep generative modelling could refine the qMRI generator, while translating multi-modal data into qMRI maps could improve the representation of physiological variability.
@@
An important extension is to explore whether the benefits of qMRI-based synthesis extend to other domains where synthetic data has been impactful, such as healthy brain parcellation [Billot2023], and super-resolution [Iglesias2023].